OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. B(0, 0), C(4, -4)

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

(0+4)/(0-4)=-1 y=x-8 x-y=8

OpenStudy (anonymous):

thnx